OpenStudy (anonymous):

Simplify each radical expression. Leave in radical form. Show your work. sqrt75 +sqrt3

OpenStudy (amorfide):

to simplify the surd you want to find the largest square number that goes into 75 this would be 25 rewrite 75 as 25x3 this gives \[\sqrt{25}\sqrt{3} +\sqrt{3}\] simplify and add together

OpenStudy (triciaal):